# Computational Analysis of Character Development in Holocaust Testimonies

### v1

@@ 6 Trajectory Clustering
Figure 4: Religious Trajectory structure distributions, from left to right: 65% of the practice trajectories have an oscillating structure and 30% are constant-active. The belief trajectories have a similar distribution, with the majority sharing a constant-positive structure (63%) and the rest distributed among oscillating (14%), constant-negative (11%), descending (6%), and ascending (5%). For the intersection of the two aspects, the two large groups cover 61% of the intersection, all have a constant-positive belief structure, while the practice valence varies: 37% oscillate and 24% have a constant-active structure.
After extracting the trajectories, we cluster them in two ways: according to a predefined taxonomy and using unsupervised methods.
We expect to identify a few main clusters, each representing a different trajectory type that can be identified by generating a path that is more similar to each of the sequences within its cluster than to any of the other sequences.
@@ 6.2 Unsupervised Hierarchical Clustering
For hierarchical clustering, we use Dynamic Time Warping Berndt and Clifford (1994) with multiple window sizes to measure the distances between the trajectories, and then run Agglomerative Zepeda-Mendoza and Resendis-Antonio (2013) and HDBSCAN clustering McInnes et al. (2017). Further details are in Appendix H.
@@ 7 Results
Besides the evaluation of the full dataset, we randomly sampled 101 testimonies and compared their predicted trajectories in three settings: prompting GPT-4o and the trained Mistral-7B on religious content predictions, and separately prompting Mistral on the full set of segments. Evaluation results for this subset are provided in Appendix G. We observe that prompting on all segments, as opposed to just religious content, leads to a significant over-prediction rate for each class. Specifically, when using the non-filtered data, the proportion of positive classes does not align with the proportion of religious content in the full dataset, as predicted by a reliable classifier trained on this data.
@@ B P P- B P P+ B+ B-
Predicted 12 12 41 7 9 10 13 127
Original scatter 22 104 48 16 71 54 14 139
Edges & middle 27 33 57 20 30 31 20 139
Equal scatter 35 24 53 23 18 25 14 180
Normal-original 24 62 51 21 52 60 20 141
@@
# predicted points 4,500 21,193 1,719 4,500 21,193 17,462 2,339 427
Table 3: Sum of $min\_sum\_dist$ for fine-tuned Mistral-7B predictions and baseline trajectories on the full dataset. Original: Sample from the distribution of the predictions. Edges&middle: Random sample from three equal splits of (0, 1); according to the predicted distribution of the label. Equal: Even scatter. Normal-original: Sample from the Gaussian with the variance and mean values of the predictions.
@@ 7.1 Clustering
Among the trajectories extracted by the pipeline, 68% contain at least two points for both practice and belief. The taxonomy reveals that, for the belief aspect, more than half (63%) share a constant positive shape, 14% oscillate, and the rest are distributed among constant negative (11%), descending (6%), and ascending (5%). For the positive trajectories, 80% cover a substantial portion of the plot, whereas negative trajectories are sparser, 59% covering up to two-thirds of the narrative. Practice trajectories have two large structures, 65% oscillate and 30% follow a constant-active pattern. 96% of the two main structures have high coverage levels. Examples for clusters are given in Appendix H.
When comparing the trends according to the two aspects, the two trends that we recognize, covering 61% of the intersection, all have a constant-positive belief structure, while the practice valence varies: 37% oscillate and 24% have a constant-active structure. This discrepancy may reflect external circumstances the survivors went through that prevented them from practicing Judaism, or social changes. These distributions are shown in Figure 4.
@@ 7.2 Clustering Evaluation
Annotators reviewed text segments from 121 trajectories and performed two tasks – classifying them according to the taxonomy and identifying the more similar pairs from a list of triplets. IAA levels for classification were substantial for belief ($\alpha=0.66$) and moderate for practice ($\alpha=0.56$), while for the triplet task, IAA was moderate ($\alpha=0.58$ for belief, $\alpha=0.32$ for practice).
@@
The DTW distances correlate with human judgment for similarity, validating its use for identifying similar trajectories. For similar pairs, the average distances are 0.46 (belief, $\sigma=0.19$) and 0.41 (practice, $\sigma=0.15$). For dissimilar pairs, these distances increase to 0.57 (belief, $\sigma=0.18$) and 0.48 (practice, $\sigma=0.20$) on average, with a t-test revealing statistical significance (belief, $p=0.01$, practice $p=0.05$).
Taxonomy annotation shows that the DTW distance also correlates with the taxonomy’s distinctions. The average distances are 0.29 (belief, $\sigma=0.12$) and 0.30 (practice, $\sigma=0.13$ for pairs of trajectories from the same class; while for pairs with different structures, the average distance is 0.32 for both aspects (belief, $\sigma=0.12$, practice, $\sigma=0.13$), with t-tests confirming statistical significance ($p<10^{-7}$ in both cases). This provides further validation for DTW by showing that part of the distance it captures is related to fundamental characteristics of the trajectory.
@@
Comparing the automatic and manual classification of trajectories into taxonomy classes, we observe F1 scores of 0.44 for both belief and practice, showing that the task is challenging, even for humans. Some discrepancies arise from the assumption that the predicted trajectories depend on the religious trajectory progressing along the narrative timeline; however, it does not hold for each case.
As with manual taxonomy classification, automatic taxonomy prediction results correlate with DTW distances (Figure 6). Pairs with the same structure exhibit smaller distances, with average distances of 0.10 (belief, $\sigma=0.04$) and 0.08 (practice, $\sigma=0.04$) for structures of the same class, compared to 0.16 (belief, $\sigma=0.05$) and 0.12 (practice, $\sigma=0.05$) for other pairs, reinforcing the link between structural alignment and DTW metrics.
@@ 8 Discussion
Our analysis identifies common structures of religious trajectories, grouping those with similar structures. For belief, the majority presents a constant positive structure, as to practice, while for practice most oscillate.
The distribution of religious content throughout the testimonies demonstrates a common narrative sequence within the dataset that justifies the comparision of trajectories on the narrative timeline (Figure 5). This can be caused by the SF interviewers’ guidelines regarding the themes and order of the testimony, including religious belief and activity.
@@
Another perspective expected to contribute to a more sophisticated picture of the patterns is including the non-positive-or-negative parts of the trajectory in the structure taxonomy analysis. The importance of incorporating these segments arises from the unsupervised clustering that identifies a separation between trajectories with multiple non-positive or negative segments, suggesting that these play a separate role in the trajectories.
Additionally, inspecting the relationship between the practice and belief trajectories, investigating whether one follows the other’s pattern or is a shift of the other, and whether one can be predicted based on the other is an idea to explore.
@@ 9 Conclusion
As for the framework’s limitations, it is important to consider the existence of a reporting bias. Our data is limited to interview transcripts, which capture the survivors’ descriptions of their beliefs without offering a broader understanding of their theological positions. Although the SF recorded testimonies in several languages and countries, it is important to note that our investigation only covered testimonies given in English.
@@ Acknowledgments
The authors acknowledge the USC Shoah Foundation - The Institute for Visual History and Education for supporting this research. We thank Dr. Naama Seri-Levi, and Dr. Mali Eisenberg for their valuable insights, and our team of annotatorts for their research assistance. Grants from the Israeli Ministry of Science and Technology, the Israeli Council for Higher Education, and the Alfred Landecker Foundation supported this research.
Appendix A Religious Content Annotation Guidelines
@@
Appendix B Practice and Belief Identification and Rating - Annotation Guidelines
@@
Appendix C GPT-4o Prompt for enlarging the portion of the minority classes
@@
Appendix D Instruction Prompts
@@
Appendix E Topic Model Based Evaluation
@@
Appendix F Thesaurus Based Evaluation
@@
    ritual circumcision (bio) mohelim Jewish religious observances Jewish schools synagogue attendance mikva’ot Yahrzeit Jewish dietary laws ghetto Jewish religious observances camp Jewish religious observances refugee camp Jewish religious observances hiding-related Jewish religious observances Islamic prayers yizkor Kaddish prison Jewish religious observances Jewish mourning customs yeshivot forced labor battalion Jewish religious observances Hasidic rebbes (bio) synagogue organizations (bio) Baalei Keriah (bio) Baalei Tefillah (bio) synagogues’ sisterhood (bio) synagogues’ men’s clubs deportation Jewish religious observances Jewish Theological Seminary of America transfer Jewish religious observances Beth Jacob schools b’nai mitzvah b’nai mitzvah (stills) Borerim Jewish religious observances (stills) Institute of Jewish Studies shamas Jewish Institute of Religion Mitnagdim observant/practicing Ramah Camping Movement Shema Yisrael
@@
    Christian religious observances church attendance religious identity communions (stills) confirmations (stills) Jehovah’s Witness missionary activities Jehovah’s Witness religious observances Mormon missionary activities Jehovah’s Witness religious beliefs camp Jehovah’s Witness religious observances baptisms forced labor battalion Jehovah’s Witness religious observances camp Christian religious observances Eucharist Islamic religious observances baptisms (stills) prison Jehovah’s Witness religious observances ghetto Christian religious observances Christian religious observances (stills) confirmations Islamic dietary laws Seventh-Day Adventist missionary activities non-observant/non-practicing Buddhist religious observances Buddhist lunar days karma Christian missionary activities Christian prayers
@@
    prayers Jewish prayers camp Jewish prayers Jewish religious beliefs prison Jewish prayers forced labor battalion Jewish prayers camp prayers ghetto prayers deportation Jewish prayers hiding-related prayers forced march Jewish prayers
@@
    Christian religious beliefs camp Jehovah’s Witness prayers Jehovah’s Witness prayers Islamic identity Buddhist religious beliefs Islamic religious beliefs Armenian Genocide faith issues Bosnian War and Genocide faith issues Guatemalan Genocide faith issues Holocaust faith issues Rwandan Tutsi Genocide faith issues
Appendix G Evaluation on a Subset of Trajectories
Evaluating a subset of the predicted trajectories presents similar trends to the evaluation in the complete dataset, results in Tables 5, 6 and 7.
@@ B P P- B P P+ B+ B-
Table 5: Sum of $min\_sum\_dist$ for GPT-4o on a subset of trajectories
@@
Table 6: Sum of $min\_sum\_dist$ of fine-tuned Mistral7B on a subset of trajectories, prompting all content
@@
Table 7: Sum of $min\_sum\_dist$ of fine-tuned Mistral7B on a subset of trajectories, prompting filtered content
Appendix H Hierarchical Clustering
@@
An example from running HDBSCAN on the belief trajectories: 16
We first truncated the values of the position points of the trajectories to two decimal points, then calulated their distance matrix using dtw_ndim.distance_matrix from the dtaidistance Python module, with window=7 and the default values for the rest of the hyper-parapeters; followed by calling HDBSCAN from the hdbscan module, with: min_cluster_size=30, min_samples=1, cluster_selection_epsilon=1 and alpha=1.
@@
An example from running HDBSCAN on the practice trajectories: 10
@@
Figure 7: (a) Active - Med
Figure 8: (b) Active
Figure 9: (d) Oscillating
@@
Figure 11: (a) Neutral
Figure 12: (b) Ascending from neutral
Figure 13: (d) Positive
Figure 14: (e) Descending Positive Neutral
Figure 15: (g) Negative to Neutral
Figure 16: Belief cluster examples

### v5

@@ 6 Trajectory Clustering
Figure 4: Religious Trajectory structure distributions, from left to right: 68% of the practice trajectories have an oscillating structure and 25% are constant-active. The belief trajectories have a similar distribution, with the majority of trajectories sharing a constant-positive structure (45%) and the rest distributed among oscillating (26%), ascending (12%), constant-negative (9%), and descending (8%). For the intersection of the two aspects, the two large groups cover 44% of the intersection, all have an oscillating practice structure, while the belief valence distributes evenly between oscillating and constant-positive structures.
After extracting the trajectories, we cluster them in two ways: according to a predefined taxonomy and using unsupervised methods. We expect to identify a few main clusters, each representing a different trajectory type that can be identified by generating a path that is more similar to each of the sequences within its cluster than to any of the other sequences.
@@ 6.2 Unsupervised Hierarchical Clustering
For hierarchical clustering, we use Dynamic Time Warping Berndt and Clifford (1994) with multiple window sizes to measure the distances between the trajectories, and then run Agglomerative Zepeda-Mendoza and Resendis-Antonio (2013) and HDBSCAN clustering McInnes et al. (2017).
@@ 7 Results
Besides the evaluation of the trajectories produced by promoting only religious content predictions, we randomly sample 101 testimonies, which we prompt and evaluate on the full set of segments. We observe that prompting on all segments, as opposed to just religious content, leads to a significant over-prediction rate for each class. See Appendix J for the full analysis.
@@ B P P- B P P+ B+ B-
Predicted 5 10 47 5 8 11 9 90
Original scatter 33 40 47 29 34 50 11 105
Edges & middle 19 29 59 19 28 29 16 99
Equal scatter 21 20 60 19 17 21 9 142
Normal-original 16 57 52 18 47 53 17 104
@@
# predicted points 6,051 23,274 1,987 6,051 23,274 20,086 3,204 917
Table 3: Sum of min_sum_distmin\_sum\_dist for fine-tuned Mistral-7B predictions and baseline trajectories on the full dataset. Original: Sample from the distribution of the predictions. Edges&middle: Random sample from three equal splits of (0, 1); according to the predicted distribution of the label. Equal: Even scatter. Normal-original: Sample from the Gaussian with the variance and mean values of the predictions.
@@ 7.1 Clustering
Among the trajectories extracted by the pipeline, 83% contain at least two points for both practice and belief. The taxonomy reveals that, for the belief aspect, 45% share a constant positive shape, 26% oscillate, and the rest are distributed among ascending (12%), constant negative (9%), and descending (8%). For the positive trajectories, 58% cover a substantial portion of the plot, whereas negative trajectories are sparser, 49% covering up to two-thirds of the narrative. Practice trajectories have two large structures, 68% oscillate and 25% follow a constant-active pattern. Over 82% of the two main structures have high coverage levels. Examples for unsupervised clusters are given in Appendix L.
When comparing the trends according to the two aspects, the three most common combinations of structures are: Positive belief & Practice Oscillating (26%), both aspects Oscillating (22%), and Positive belief & Active Practice (17%). These distributions are shown in Figure 4, further analysis in Appendix K.
@@ 7.2 Clustering Evaluation
Annotators reviewed text segments from 121 trajectories and performed two tasks – classifying them according to the taxonomy and identifying the more similar pairs from a list of triplets. IAA levels for classification were substantial for belief ($\alpha=0.66$) and moderate for practice ($\alpha=0.56$), while for the triplet task, IAA was moderate ($\alpha=0.58$ for belief, $\alpha=0.32$ for practice), showing that the task is challenging, even for humans.
@@
The DTW distances correlate with human judgment for similarity, validating its use for identifying similar trajectories. For similar pairs in the belief aspect, the average distance is 0.42 ($\sigma=0.19$), whereas for dissimilar pairs, it increases to 0.53 ($\sigma=0.18$), with a t-test revealing statistical significance ($p<0.006$). However, this pattern was not observed for the practice aspect.
Taxonomy annotation shows that the DTW distance also correlates with the taxonomy’s distinctions. The average distances are 0.38 (belief, $\sigma=0.14$) and 0.41 (practice, $\sigma=0.16$) for pairs of trajectories from the same class; while for pairs with different structures, the average distance is 0.43 (belief, $\sigma=0.14$), 0.44 (practice, $\sigma=0.16$). Running a t-test confirms statistical significance ($p<10^{-12}$ in both cases). This provides further validation for DTW by showing that part of the distance it captures is related to fundamental characteristics of the trajectory.
@@
Comparing the automatic and manual classification of trajectories into taxonomy classes, we observe F1 scores of 0.40 (belief) and 0.38 (practice) for both belief and practice. Some discrepancies arise from the assumption that the predicted trajectories depend on the religious trajectory progressing along the narrative timeline; however, this does not hold for each individual case.
As with manual taxonomy classification, automatic taxonomy prediction results correlate with DTW distances (Figure 6). Pairs with the same structure exhibit smaller distances, with average distances of 0.11 (belief, $\sigma=0.05$) and 0.10 (practice, $\sigma=0.04$) for structures of the same class, compared to 0.16 (belief, $\sigma=0.06$) and 0.12 (practice, $\sigma=0.04$) for other pairs, reinforcing the link between structural alignment and DTW metrics.
@@ 8 Discussion
Our analysis demonstrates the value of this test case by identifying common structures of religious trajectories and grouping those with similar structures. For belief, the majority of trajectories present a constant positive structure, whereas for practice, most trajectories oscillate.
The distribution of religious content throughout the testimonies, demonstrating a common narrative sequence within the dataset, lends further support to the comparison of trajectories on the narrative timeline (Figure 5). This can be caused by the SF interviewers’ guidelines, which provide instructions for conducting the interviews. [^9] http://tiny.cc/sf-interview-guidelines
While these patterns reveal commonalities, interpreting them requires caution due to potential model biases. The confusion matrices (see Appendix M) reveal that 64% of belief statements with gold label Other were falsely predicted as Positive, and 56% of practice–Other were misclassified as Active. This aligns with the idea of positivity bias (Ashwin et al., 2023; Fatahi et al., 2024), where models tend to overpredict positive categories. In addition, because this work engages with the sensitive domain of religion, it is important to acknowledge potential stereotype biases rooted in pretraining data, which may introduce unintended sentiments Elsafoury and Abercrombie (2023); Garcia and Weilbach (2023). For example, a Christian prayer (e.g., addressed to “Dear Lord” or “Mother Maria”) is often predicted with a positive valence, even though our annotation guidelines define any non-Jewish religious expressions as carrying a negative valence.
@@
Another perspective expected to contribute to a more nuanced picture of the patterns would include the non-positive-or-negative parts of the trajectory in the structure taxonomy analysis. The importance of incorporating these segments arises from the unsupervised clustering that induces a distinction between trajectories with multiple non-positive or negative segments, suggesting that these play a separate role in the trajectories.
Additionally, inspecting the relationship between practice and belief trajectories, investigating whether one follows the other’s pattern or is a shift of the other, and whether one can be predicted based on the other is an interesting avenue for future work.
@@ 9 Conclusion
As for the framework’s limitations, it is important to consider the existence of reporting and survivorship bias in existing historical texts Chen et al. (2024). Our data is limited to the interview transcripts, which capture the survivors’ descriptions of their beliefs told in hindsight, without offering a broader understanding of their theological positions. Although the SF recorded testimonies in several languages and countries, it is important to note that our investigation only covered testimonies given in English.
@@ Acknowledgments
The authors acknowledge the USC Shoah Foundation – The Institute for Visual History and Education for supporting this research. We thank Dr. Naama Seri-Levi, and Dr. Mali Eisenberg for their valuable insights, and our team of annotators for their research assistance. Grants from the Israeli Ministry of Science and Technology, the Israeli Council for Higher Education, and the Alfred Landecker Foundation supported this research.
Appendix A The relationship between the chronological order of the testimonies and the narrative timeline
Wagner et al. (2022) mapped the SF thesaurus keywords into periods of: before, during, and after the war, and reflection descriptions. For each segment in the data that matches one of these periods, we recorded its average position on the narrative timeline and got the pattern of: [before: 0.35, during: 0.45, after: 0.55, reflection: 0.71]. Both the distribution of the segments that match these keywords along the narrative timeline and the distribution of the religious content validate the comparison of the trajectories along the narrative timeline (see Figure 7).
The SF interviewers’ guidelines include discussing pre-war religious activities at the beginning, and the survivor’s present views toward the end, which strengthens this observation. In addition, the lengths of the responses are longer in the middle than at the beginning and the end, suggesting that the middle of the testimonies contains stories that the witness accounts in his choice of order.
Regarding the structure of belief trajectories, particularly the meaning of oscillations, while belief descriptions can convey meaning about how belief is talked about, many of the belief segments reflect past beliefs, for example, “Every night I would take my prayer book… and cry”. However, there is always tension between one’s personal life as reflected in conversation in the present and their historically lived experience Rosenthal (2006). In this analysis, we concentrate on the former.
Figure 7: Distributions of the segments that match keywords mapped to periods of: before, during, and after the war, and reflection descriptions, along the narrative timeline
Appendix B Religious Content Annotation Guidelines
@@
Appendix C Practice and Belief Identification and Rating - Annotation Guidelines
@@
Appendix D Annotations
Table 4: The distribution of all the annotations collected.
Table 5: The label distribution for the overlap subset of annotations.
Annotation Task Class Annotations Collected
Religious content identification Positive 674
Negative 4321
Practice valence classification Active 343
Inactive 77
Other-practice 49
None 607
Belief valence classification Positive 122
Negative 68
Other-belief 44
None 862
Table 4: Breakdown of annotations collected across three annotation tasks: (1) Religious content identification, (2) Practice valence classification, and (3) Belief valence classification.
Annotation Task Label Distribution
(Agreed Upon Samples)
Agreement Size Overlap Size
Religious content classification Positive: 227
Negative: 520
747 833 (21%)
Practice valence Active: 84
Inactive: 40
Other: 32
None: 189
345 540 (46%)
Belief valence Positive: 37
Negative: 34
Other: 34
None: 314
417 540 (46%)
Table 5: Overlap label distribution for both annotation tasks. Agreement size is the number of pairs agreed open; Overlap size is the total number of overlapping samples between annotators.
Appendix E GPT-4o Prompt for enlarging the portion of the minority classes
@@
Appendix F Instruction Prompts
@@
Appendix G Model Selection F1 scores
Table 6: A summary of the performance across models on the test data for rating practice and belief valence.
Model Prompt ID Practice Belief
GPT-4o few-shot 0.51 0.60
Fine-tuned Mistral zero-shot 0.43 0.45
GPT-4o zero-shot 0.47 0.55
Mistral zero-shot 0.26 0.25
GPT-4.1 few-shot 0.55 0.64
GPT-4.1-mini few-shot 0.48 0.58
GPT-4.1-nano few-shot 0.27 0.33
GPT-4.1 zero-shot 0.48 0.62
GPT-4.1-mini zero-shot 0.48 0.55
GPT-4.1 1-shot 0.50 0.62
Table 6: Macro F1 scores for the various models and settings tested for selecting models to produce trajectories with
Appendix H Topic Model Based Evaluation
@@
Appendix I Thesaurus Based Evaluation
@@
    Ritual circumcision · (bio) mohelim · Jewish religious observances · Jewish schools · synagogue attendance · mikva’ot · Yahrzeit · Jewish dietary laws · ghetto Jewish religious observances · camp Jewish religious observances · refugee camp Jewish religious observances · hiding-related Jewish religious observances · Islamic prayers · yizkor · Kaddish · prison Jewish religious observances · Jewish mourning customs · yeshivot · forced labor battalion Jewish religious observances · Hasidic rebbes · (bio) synagogue organizations · (bio) Baalei Keriah · (bio) Baalei Tefillah · (bio) synagogues’ sisterhood · (bio) synagogues’ men’s clubs · deportation Jewish religious observances · Jewish Theological Seminary of America · transfer Jewish religious observances · Beth Jacob schools · b’nai mitzvah · b’nai mitzvah (stills) · Borerim · Jewish religious observances (stills) · Institute of Jewish Studies · shamas · Jewish Institute of Religion · Mitnagdim · observant/practicing · Ramah Camping · Movement · Shema Yisrael.
@@
    Christian religious observances · church attendance · religious identity · communions (stills) · confirmations (stills) · Jehovah’s Witness missionary activities · Jehovah’s Witness religious observances · Mormon missionary activities · Jehovah’s Witness religious beliefs · camp Jehovah’s Witness religious observances · baptisms · forced labor battalion Jehovah’s Witness religious observances · camp Christian religious observances · Eucharist · Islamic religious observances · baptisms (stills) · prison Jehovah’s Witness religious observances · ghetto Christian religious observances · Christian religious observances (stills) · confirmations · Islamic dietary laws · Seventh-Day Adventist missionary activities · non-observant/non-practicing · Buddhist religious observances · Buddhist lunar days · karma · Christian missionary activities · Christian prayers.
@@
    Prayers · Jewish prayers · camp Jewish prayers · Jewish religious beliefs · prison Jewish prayers · forced labor battalion Jewish prayers · camp prayers · ghetto prayers · deportation Jewish prayers · hiding-related prayers · forced march Jewish prayers.
@@
    Christian religious beliefs · camp Jehovah’s Witness prayers · Jehovah’s Witness prayers · Islamic identity · Buddhist religious beliefs · Islamic religious beliefs · Armenian Genocide faith issues · Bosnian War and Genocide faith issues · Guatemalan Genocide faith issues · Holocaust faith issues · Rwandan Tutsi Genocide faith issues.
Appendix J Evaluation on a Subset of Trajectories
Besides the evaluation of the full dataset, we randomly sampled 101 testimonies and compared their predicted trajectories in three settings: prompting GPT-4o and the trained Mistral-7B on religious content predictions, and separately prompting Mistral on the full set of segments. Evaluation results for this subset are provided in Tables 8, 9 and 10. We observe that prompting on all segments, as opposed to just religious content, leads to a significant over-prediction rate for each class. Specifically, when using the non-filtered data, the proportion of positive classes does not align with the proportion of religious content in the full dataset, as predicted by a reliable classifier trained on this data.
@@ B P P- B P P+ B+ B-
Table 8: Sum of min_sum_distmin\_sum\_dist for GPT-4o on a subset of trajectories
@@
Table 9: Sum of min_sum_distmin\_sum\_dist of fine-tuned Mistral7B on a subset of trajectories, prompting all content
@@
Table 10: Sum of min_sum_distmin\_sum\_dist of fine-tuned Mistral7B on a subset of trajectories, prompting filtered content
Appendix K Structure-Based Taxonomy Analysis
To examine how density affects the structural distribution of the trajectories, we plot the structure distribution for multiple bins, the sizes of which are inspired by the trajectory length distribution. Results in Figures: 8, 9, 10.
Figure 8: Religious Trajectory structure distributions, for 134 trajectories of length 2-3 for belief and 2-13 for practice.
Figure 9: Religious Trajectory structure distributions, for 215 trajectories of length 4-8 for belief and 14-29 for practice.
Figure 10: Religious Trajectory structure distributions, for 131 trajectories of length > 8 for belief and > 29 for practice.
Appendix L Hierarchical Clustering
@@
An example from running HDBSCAN on the belief trajectories: Figure 11
We first truncated the values of the position points of the trajectories to two decimal points, then calculated their distance matrix using dtw_ndim.distance_matrix from the dtaidistance Python module, with window=7 and the default values for the rest of the hyperparameters; followed by calling HDBSCAN from the hdbscan module, with: min_cluster_size=30, min_samples=1, cluster_selection_epsilon=1 and alpha=1.
@@
An example from running HDBSCAN on the practice trajectories: Figure 12
@@
(a) Other, Ascending
(b) Positive, Other
(c) Other and Positive oscillating
(d) Positive
(e) Positive, Other, Positive
(f) Sparse Negative
Figure 11: Belief cluster examples
(a) Active, Other, Active
(b) Active
(c) Oscillating
(d) Active, Inactive, Active
@@
Appendix M Classification Performance Analysis
The confusion matrices in Figure 13 summarize the classification performance.
Figure 13: Practice and Belief classification confusion matrices